OpenStudy (anonymous):

Which fractions are equivalent to -15/20?

OpenStudy (gabylovesyou):

Some equivalent fractions of -15/20 are: -3/4 = -6/8 = -9/12 = -12/16 = -15/20 = -18/24 = -21/28 = -24/32 = -27/36 = -30/40 = -33/44 = -36/48 = -39/52 = -42/56 = -45/60 = -48/64 = -51/68 = -54/72 = -57/76 = -60/80 http://www.calculatorsoup.com/calculators/math/fractionsequivalent.php

OpenStudy (anonymous):

Awesome! thank you so much!

OpenStudy (gabylovesyou):

no prob :D

OpenStudy (anonymous):

\(\bbox[5pt,skyblue ,border:5px solid blue]{I <3 ~LaTeX}\)